# ScholaWrite: A Dataset of End-to-End Scholarly Writing Process

### v2

@@ Abstract
Writing is a cognitively demanding task involving continuous decision-making, heavy use of working memory, and frequent switching between multiple activities. Scholarly writing is particularly complex as it requires authors to coordinate many pieces of multiform knowledge. To fully understand writers’ cognitive thought process, one should fully decode the end-to-end writing data (from individual ideas to final manuscript) and understand their complex cognitive mechanisms in scholarly writing. We introduce ScholaWrite dataset, the first-of-its-kind keystroke logs of an end-to-end scholarly writing process for complete manuscripts, with thorough annotations of cognitive writing intentions behind each keystroke. Our dataset includes LaTeX-based keystroke data from five preprints with nearly 62K total text changes and annotations across 4 months of paper writing. ScholaWrite shows promising usability and applications (e.g., iterative self-writing) for the future development of AI writing assistants for academic research, which necessitate complex methods beyond LLM prompting. Our experiments clearly demonstrated the importance of collection of end-to-end writing data, rather than the final manuscript, for the development of future writing assistants to support the cognitive thinking process of scientists. Our de-identified dataset, demo, and code repository are available on our project page111https://minnesotanlp.github.io/scholawrite/.
@@ 1 Introduction
Although much literature has examined the cognitive processes used in general writing tasks Lindgren and Sullivan (2003); Hayes and Berninger (2014); Zhu et al. (2023); Wengelin and Johansson (2023), the specific cognitive patterns in scholarly writing remain understudied. Our work observes the end-to-end writing process of scholarly writing, inspired by the method of keystroke collections that have been a major methodology to observe individual writing processes Baaijen et al. (2012); Latif (2008); Leijten and Van Waes (2013) in cognitive science or a few recent works in NLP at a small scale Koo et al. (2023); Velentzas et al. (2024).
We believe our work (ScholaWrite) is the first attempt to present a keystroke corpus of scholarly writing with annotations of cognitive writing processes, which were collected over multiple months and produced by early-career researchers. Furthermore, we developed novel data collection and annotation systems using Chrome extension, which improves data accessibility to the public and enables real-time collection of LaTeX-based keystrokes from multiple researchers. We also present a comprehensive taxonomy of cognitive writing processes specific to the scholarly writing domain. Throughout several experiments with LLMs, ScholaWrite shows promising usability as a resource for enabling LLMs to understand the human cognitive process of scholarly writing and provide writing suggestions that are aware of writers’ cognitive behaviors.
@@
We present ScholaWrite, a curated dataset of nearly 62K LaTeX-based keystrokes that were turned into publications in the computer science domain, annotated by experts in linguistics and computer science. We develop a taxonomy of scholarly writing intentions, providing an overall understanding of how scholars tend to develop writing manuscripts. Experiment results showed that the Llama-8B model fine-tuned on the ScholaWrite dataset achieved the high linguistic quality of the final writing.
@@ 2.1 Cognitive Theory of Writing Process
Research on human writing significantly shifted from focusing solely on the final written product to exploring the cognitive stages of various writing phases Diederich (1974); Perl (1979); Krapels (1990); MacArthur and Graham (2016).
Building on this process-oriented approach, Flower and Hayes (1981)’s cognitive writing theory outlines three key cognitive sub-processes grounded in empirical evidence: (1) planning; (2) translating; and (3) reviewing. These processes are complex, interconnected, and non-linear, with writers moving fluidly between them. Our work expands this model to develop our own taxonomy of more granular cognitive patterns of human writing, specifically in scholarly communication.
Koo et al. (2023) presented a taxonomy of the scholarly writing process based on keystroke collections of just a short research plan over a thirty-minute session. Inspired by this work, we extend their approach to a larger-scale collection of keystrokes over months, whose final written products turned into research publications. Reviewed by a linguistics expert, our comprehensive taxonomy is built on top of prior literature and encompasses end-to-end writing trajectories of the cognitive process in a scholarly domain.
@@ 2.2 Keystroke Loggers for Scholarly Writing
Keystroke logging tools such as Inputlog allow researchers to observe digital writing without interfering with the writer’s process Chan (2017); Johansson et al. (2010); Leijten and Van Waes (2013); Lindgren and Sullivan (2019). However, they tend to operate in a closed ecosystem, i.e., only running in MS Word. This makes the current tools less accessible, especially for the scientific communities who use other typesetting such as LaTeX. Moreover, it is not well-suited for collecting data from extended writing sessions.
Currently, such keystroke logging studies have rarely examined scholarly communication settings that use LaTeX, especially for scientific writing purposes. To address this gap, our work implemented novel several interconnected systems that securely collect keystrokes in real-time over months from Overleaf, a widely-used online LaTeX editor, store them without privacy issues, and retrieve them for annotating writing intentions at span level. Our system workflow provides a more natural environment for observing cognitive writing activities by multiple researchers over long periods and across multiple sections without interrupt their writing.
@@ 2.3 Datasets for Scholarly Writing
Previous work primarily focused on constructing datasets for understanding the writing process of scholarly domains, which vary in their content and scale. Much of the publicly available datasets analyzed linguistic style changes or grammatical edits during the revision process Du et al. (2022b); Jiang et al. (2022); Ito et al. (2019); Mita et al. (2022). Also, the datasets from D’Arcy et al. (2024); Jourdan et al. (2024); Kuznetsov et al. (2022) collected edits conditioned on feedback and peer review process, while Kobayashi et al. (2022); Narimatsu et al. (2021) constructed datasets for citation generations. Existing work is also limited to only specific sections of scientific papers (e.g., abstract, introduction) Du et al. (2022b); Mita et al. (2022). In comparison, we construct a dataset that covers all cognitive phases of the writing process from its beginning to the final written product, through long periods of the data collection process.
Furthermore, those existing corpora for scientific writing research are created from the comparison between multiple versions of full-length, final products retrieved from preprint databases Jiang et al. (2022); Du et al. (2022b); D’Arcy et al. (2024); Jourdan et al. (2024). As pointed out by Jourdan et al. (2023), the comparison of finalized outputs does not present how those manuscripts evolved. Our work addresses this problem by building a corpus of keystrokes each of which represents the progress of the publications (See Figure 2).
@@ 3 Dataset Construction
To preserve writers’ natural thought flow and collect authentic writing data, we developed a Chrome extension that records keystroke data in the background. After the study, an in-depth annotation process is conducted to interpret the intent behind each keystroke. For open research, we make all the code of our data collection and annotation framework publicly available: https://minnesotanlp.github.io/scholawrite/. We describe the detailed process of our data collection, annotation system design and taxonomy creation in the following sections.
@@ 3.1 Participant Recruitment
We recruited 10 graduate students222All of them currently attend a four-year university in the U.S. and are proficient in English. from the computer science department who were actively preparing academic manuscripts in Overleaf LaTeX editor333https://overleaf.com/ for submission to peer-reviewed conferences in AI and NLP. The data is collected from November 2023 to February 2024, totaling up to 4 months. Our study has been approved by the IRB institution of the authors. Please refer to Appendix A.1 for more detailed descriptions.
@@ 3.2 Chrome Extension for Keystroke Collection in Overleaf
We designed and implemented a Chrome extension (Figure 4(a)(a)), which enables the real-time collection of keystroke trajectories in the Overleaf platform. Participants can create their account credentials, and after logging into the system, the extension monitors the user’s keystrokes in the background silently, without disrupting the typical writing process.
When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel. When each of these actions444Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page. occurs, the extension uses ‘diff_match_patch’ package555https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 3). Then, the extension will send the difference array along with other metadata (e.g., time stamp, author ID, file name, etc.) to the backend server. To prevent any issue of private data collection, the backend fetches only the IDs of the Overleaf projects that participants consented to share during the recruitment process and filters out participants’ keystroke data from any unauthorized projects666We used Google Sheet API to retrieve ID information that we collected during the recruitment process..
For any Overleaf project that consists of multiple LaTex files, which is common in collaborative scholarly writing, we also collected all keystrokes from subfiles associated with the main LaTeX file. Our comprehensive data collection process captures the end-to-end writing processes of the participating authors across all parts of Overleaf projects. This approach ensures that our dataset reflects the full scope of scholarly writing including edits made in auxiliary files such as files of each section, appendix, bibliography, etc.
All keystroke data stored in the database were processed for annotation and model training purposes. Here, raw keystrokes from the Chrome extension system are stored as arrays that represent the differences between subsequent keystrokes, along with accompanying metadata.
@@
| Planning | Idea Generation | Formulate and record initial thoughts and concepts. | writing down keywords of a paragraph beforehand (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
@@ 3.3 Writing Intention Annotation
To capture comprehensive writing behaviors, we designed a data collection process that involves running a Chrome extension in the background. However, keystroke data alone may be insufficient to fully understand the nuanced intentions behind each writing action. To bridge this gap, we implement an additional annotation step to decode the writer’s intent for each keystroke using a specialized annotation interface. This process is crucial for developing a detailed taxonomy of the end-to-end scholarly writing process, which serves as the basis for annotating the collected keystrokes. Further details about the annotation process can be found in Appendix B.
To perform annotations, we first processed those raw keystroke entries by file name, type of writing actions (e.g., add, delete), an array of differences between two subsequent texts, and line number in an Overleaf editor.
@@
We developed a novel interface for the keystroke annotation process. The interface offers various modes for visualizing keystroke collections within each Overleaf project: by time, LaTeX file, and author (Figure 4(a)(b)).
In the right panel, annotators can interactively right-click to navigate the timeline of collected keystrokes in different colors (e.g., red for deletions and green for additions). They then identify a range of keystrokes whose underlying purpose or goal is similar when making edits during the writing process (e.g., generating outlines or correcting grammatical errors). They can manually enter the beginning and ending indices and annotate the span by selecting a pre-defined label of the relevant intention from the dropdown menu in the left panel. Furthermore, the interface shows the metadata information of each keystroke, along with the selected label.
@@
Following Pustejovsky et al. (2017)’s iterative annotate-and-refine framework, we performed an iterative open coding for the collected keystrokes from the two Overleaf projects. Using the annotation interface, each annotator reviewed the first few writing keystrokes and identified the high-level processes based on (Flower and Hayes, 1981)’s initial taxonomy of cognitive writing (planning, implementation, and revision). Next, the annotators (1) determine the span of keystrokes that form a meaningful unit together (e.g., a clause or sentence) within the high-level process and (2) decide the underlying intention of edits. The annotators coded these spans independently, resolving conflicts through multiple in-depth discussions. Finally, a cognitive linguist participated in all discussions, refining the definition and scope of each writing intention.
To ensure the validity of our taxonomy, we followed the two standard practices Nickerson et al. (2013); Kundisch et al. (2021): (1) All elements should be mutually exclusive and (2) the taxonomy should be collectively comprehensive. To achieve this, we flagged any ambiguous keystroke spans that could be interpreted as representing multiple intentions. Through iterative discussion, annotators refined these spans’ characteristics and improved each category’s decision boundaries. Once the codebook and taxonomy were finalized, we updated intention annotations777The list of those intentions is then updated in the dropdown bar of the annotation interface. for the selected two projects and applied the taxonomy to the remaining projects. The final inter-annotator agreement is using a weighted F1 score in a multi-label, multi-class setting for a randomly selected window of 1K keystrokes. Please refer to Appendix B for more details about the entire annotation procedures.

### v3

@@ Abstract
Writing is a cognitively demanding task involving continuous decision-making, heavy use of working memory, and frequent switching between multiple activities. Scholarly writing is particularly complex as it requires authors to coordinate many pieces of multiform knowledge. To fully understand writers’ cognitive thought process, one should fully decode the end-to-end writing data (from individual ideas to final manuscript) and understand their complex cognitive mechanisms in scholarly writing. We introduce ScholaWrite dataset, a first-of-its-kind keystroke corpus of an end-to-end scholarly writing process for complete manuscripts, with thorough annotations of cognitive writing intentions behind each keystroke. Our dataset includes LaTeX-based keystroke data from five preprints with nearly 62K total text changes and annotations across 4 months of paper writing. ScholaWrite shows promising usability and applications (e.g., iterative self-writing), demonstrating the importance of collection of end-to-end writing data, rather than the final manuscript, for the development of future writing assistants to support the cognitive thinking process of scientists. Our de-identified data examples and code are available on our project page111https://minnesotanlp.github.io/scholawrite/.
@@ 1 Introduction
Although much literature has examined the cognitive processes used in general writing tasks Lindgren and Sullivan (2003); Hayes and Berninger (2014); Zhu et al. (2023); Wengelin and Johansson (2023), the specific cognitive patterns in scholarly writing remain understudied. Our work observes the end-to-end cognitive process of scholarly writing, inspired by the method of keystroke collections that have been a major methodology to observe individual writing processes Baaijen et al. (2012); Latif (2008); Leijten and Van Waes (2013) in cognitive science or a few recent works in NLP at a small scale Koo et al. (2023); Velentzas et al. (2024).
We believe our work (ScholaWrite) is the first attempt to present a keystroke corpus of scholarly writing with annotations of cognitive writing intentions, which were collected over multiple months and produced by early-career researchers. Furthermore, we developed novel data collection and annotation systems, which improves data accessibility to the public and enables real-time collection of LaTeX-based keystrokes from multiple researchers. We also present a comprehensive taxonomy of cognitive writing processes specific to the scholarly writing domain. Throughout several experiments with LLMs, ScholaWrite shows promising usability as a resource for enabling LLMs to understand the human cognitive process of scholarly writing and provide writing suggestions that are aware of writers’ cognitive behaviors.
@@
We present ScholaWrite, a curated dataset of nearly 62K LaTeX-based keystrokes that were turned into publications in the computer science domain, annotated by experts in linguistics and computer science. We develop a taxonomy of cognitive process of scholarly writing, providing an overall understanding of how scholars tend to develop writing manuscripts. Experiment results showed that the Llama-8B model fine-tuned on the ScholaWrite dataset achieved the high linguistic quality of the final writing.
@@ 2.1 Cognitive Theory of Writing Process
Research on human writing shifted from analyzing final written products to examining cognitive processes across various writing phases Diederich (1974); Krapels (1990); MacArthur and Graham (2016). Building on this process-oriented approach, Flower and Hayes (1981)’s cognitive writing theory outlines three key sub-processes: (1) planning, (2) translating, and (3) reviewing. These interconnected and non-linear stages of human writing inform our work, where we expand this model to a more granular taxonomy of cognitive writing patterns in scholarly communication.
Koo et al. (2023) introduced a taxonomy of scholarly writing process based on keystroke data from short research plans written in 30-minute sessions. Expanding on this, we perform a larger-scale keystroke collection over months, whose final products culminated in research publications. Expert-reviewed and grounded in prior literature, our taxonomy encompasses end-to-end cognitive trajectories of scholarly writing.
@@ 2.2 Keystroke Loggers for Scholarly Writing
Keystroke logging tools (e.g., Inputlog) allow researchers to observe digital writing without disrupting the writing process Chan (2017); Johansson et al. (2010); Leijten and Van Waes (2013); Lindgren and Sullivan (2019). However, they are often restricted to a closed ecosystem, such as MS Word, and thus less accessible for scientific communities who use LaTeX. Moreover, they are not well-suited for collecting data from extended writing sessions.
Current studies rarely examined scholarly communication that uses LaTeX, particularly for scientific writing. To address this gap, we developed systems that securely collect real-time keystrokes over months from Overleaf, a widely-used online LaTeX editor, while ensuring privacy. The system workflow enables span-level annotation of writing intentions, offering a natural, uninterrupted environment for studying cognitive writing activities over long periods and across multiple sections.
@@ 2.3 Datasets for Scholarly Writing
Previous work primarily focused on constructing datasets to analyze scholarly writing processes, which vary in content and scale. Publicly available datasets tend to track linguistic style changes or grammatical edits during revision Du et al. (2022b); Jiang et al. (2022); Ito et al. (2019); Mita et al. (2022), while others capture edits based on feedback and peer review D’Arcy et al. (2024); Jourdan et al. (2024); Kuznetsov et al. (2022) or citation generation Kobayashi et al. (2022); Narimatsu et al. (2021). Also, most focus on specific sections of papers (e.g., abstracts, introductions) Du et al. (2022b); Mita et al. (2022). In contrast, our dataset covers all cognitive phases of writing, from inception to final manuscript, over an extended period.
Furthermore, those existing corpora often compare multiple versions of final manuscripts from preprint databases Jiang et al. (2022); Du et al. (2022b); D’Arcy et al. (2024); Jourdan et al. (2024), missing how those manuscripts evolved Jourdan et al. (2023). To address this, we build a keystroke-based corpus that captures the progression of publications in real-time (See Figure 2).
@@ 3 Dataset Construction
To collect authentic writing data from writers’ natural thoughts, we developed a Chrome extension that records keystroke data in the background. After the study, an in-depth annotation process is conducted to interpret the intent behind each keystroke. We describe the detailed process of our data collection, annotation system design, and taxonomy creation in the following sections.
@@ 3.1 Participant Recruitment
We recruited 10 graduate students222All participants attend a R1 university in US and are proficient in English. Appendix A.1 details the recruitment process. from the computer science department who were actively preparing academic manuscripts in Overleaf LaTeX editor for submission to peer-reviewed conferences in AI and NLP. The data is collected from November 2023 to February 2024, totaling up to 4 months. Our study has been approved by the IRB institution of the authors. Please refer to Appendix A.1 for more detailed descriptions.
@@ 3.2 Chrome Extension for Keystroke Collection in Overleaf
We designed and implemented a Chrome extension (Appendix Figure 7), which enables the real-time collection of keystroke trajectories in the Overleaf platform. Participants can create their account credentials, and after logging into the system, the extension monitors the user’s keystrokes in the background silently without disrupting writing process. See Appendix A.2 for more description of the extension workflow.
@@
Planning | Idea Generation | Formulate and record initial thoughts and concepts. | writing keywords or notes (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
@@ 3.3 Writing Intention Annotation
To decode the nuanced intentions behind each writing action, we annotate each keystroke collected from the Chrome extension, using a specialized interface. This process is crucial for developing a detailed taxonomy of the end-to-end scholarly writing process, which serves as the basis for annotating the collected keystrokes. To perform annotations, we processed those raw keystroke entries by file name, type of writing actions, an array of differences between two subsequent texts, and line number in an Overleaf editor. Please see Appendix B for further annotation details.
@@
We developed a novel interface for keystroke annotation, enabling visualization by time, LaTeX file, and author (Appendix Figure 8). Annotators can navigate the keystroke timeline, identify a span of related edits (e.g., outlining or grammar correction), and assign a pre-defined intention label from the dropdown menu.
@@
Following Pustejovsky et al. (2017), we performed an iterative open coding for the collected keystrokes from the two Overleaf projects. Using the annotation interface, each annotator reviewed the first few writing keystrokes and identified the high-level processes. Next, the annotators (1) determine the span of keystrokes that form a meaningful unit together (e.g., a clause or sentence) within the high-level process and (2) decide the underlying intention of edits. The annotators coded these spans independently, resolving conflicts through multiple in-depth discussions. Finally, a cognitive linguist participated in all discussions, refining the definition and scope of each writing intention.
To ensure taxonomy validity, we followed two standard practices Nickerson et al. (2013); Kundisch et al. (2021): (1) mutual exclusivity and (2) collective comprehensiveness. Ambiguous keystroke spans were flagged and refined through iterative discussions to sharpen category boundaries. After finalizing the codebook and taxonomy, we updated intention annotations333The updated list appears in annotation interface. for the two selected projects and applied the taxonomy to the rest. The final inter-annotator agreement was (weighted F1 score in a multi-label setting for a 1K-keystroke sample444Unlike Kappa, which assumes categorical agreement and struggles with multi-label classification, weighted F-1 scores better handle class imbalances and partial agreements.). See Appendix B for full annotation details.
@@ 4 ScholaWrite: A Dataset of Cognitive Process of Scholarly Writing
Summary
Our post-processed dataset, ScholaWrite, contains end-to-end writing trajectories with annotated intentions, for five Overleaf projects whose final product turned into arXiv preprints. The dataset consists of 61,504 arrays of keystroke changes with fine-grained annotations of writing intentions. Appendix Table 6 shows the overall statistics of the dataset.
Writing Intention Distributions
The writing intention distributions is illustrated in Figure 3(c). During the planning stage, idea generation is the most frequent intention. Text production dominates the implementation stage, and clarity becomes the primary focus during the revision stage. Since we defined the boundary between implementation and revision as the point where meaning changes occur, the high frequency of text production suggests that authors often continue revising their texts by introducing new topics into the documents. Appendix Table 8 shows the label distribution per project.
Label | Subsequent label | Probability | |
Idea Generation | Text Production | 0.52 | |
Idea Organization | Idea Generation | 0.34 | |
Citation Integration | Text Production | 0.37 | |
Cross-reference | Text Production | 0.36 | |
Clarity | Text Production | 0.35 | |
Coherence | Text Production | 0.34 | |
Scientific Accuracy | Text Production | 0.34 |
Flows Between Writing Intentions
We analyze the flow of writing intentions most likely to follow each other throughout the writing process, as presented in Table 2. Many intentions proceed to text production, where the idea generation has the highest probability of feeding into text production. Figure 3(a) also shows a similar pattern of flows between text productions and other intentions. Also, in Table 2 the reflexive relationship between text production and clarity suggests a constant loop of producing new texts and refining them, ensuring the precision and reasonability of their claims. Please see Appendix Figure 13 for the intention flows of each of the five projects.
Time-series Distributions of Writing Intentions
We calculated the average Wasserstein distance between each intention distribution and uniform distribution to assess how evenly intentions occur throughout the writing process. As shown in Figure 3(d), text production had the smallest distance, while scientific accuracy had the largest. This suggests that text production is spread throughout the writing, whereas scientific accuracy is concentrated in specific phases. A similar pattern appears in Figure 3(e) , where text production is evenly distributed over time, and scientific accuracy peaks in the middle-to-late of the writing process. Figure 3(b) demonstrates a similar pattern that text production of the implementation stage spreads evenly over time. Types of clarity and structural revisions tend to occur after the medium steps of writing. Please see Appendix Figures 9 to 12 for each of the five projects.
5 Applications for Writing Assistance
We envision ScholaWrite as a valuable resource for training language models and improving future writing assistants for scholarly writing. To evaluate its usability, we conducted experiments training LLMs to mimic the complex, non-linear writing processes of human scholars. Specifically, we aimed to showcase the capabilities of LLMs trained on ScholaWrite in two scenarios:
(1) Predicting an author’s next writing intention (§5.1): The task is crucial for writing assistants to accurately assess the writer’s current status in context and predict the correct writing intention. This enables them to offer cognitively-appropriate writing suggestions that align with the writer’s needs.
(2) Iteratively generating scholarly writing actions from scratch (§5.2) (called Iterative Self-Writing), mirroring the human writing process: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
5.1 Predicting Next Writing Intention
Setup & Metrics
This task (represented by the inner box “Prediction” in Figure 4) takes the “before-text” from a keystroke pair (i.e., Listing LABEL:table:single-entry) and context prompt, and predicts the writing intention to apply for the subsequent actions.
We use BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and Llama3.1-8B-Instruct Dubey et al. (2024) as baselines, fine-tuning each on the ScholaWrite training set. For comparison, we also run GPT-4o OpenAI (2024) on the test set. To evaluate model performance, we used a weighted F-1 score555Weighted F-1 was chosen to address skewed label distribution (as shown in Figure 3(c) and Table 1).. See Appendix E.2 for details.
BERT | RoBERTa | Ll ama-8B | GPT-4o | |
Base | 0.04 | 0.02 | 0.12 | 0.08 |
+ SW | 0.64 | 0.64 | 0.13 | - |
Iter. | Llama-8B-SW | Llama-8B-Zero | GPT-4o |
[…]
Results
Table 3 presents the weighted F1 scores for predicting writing intentions across baselines and fine-tuned models. Regardless of the intricate nature of the task itself666Each model predicts the next intention using only the ”before” text, while human annotators consider multiple ”before and after” edits. Moreover, the chosen next intention is not necessarily the only correct one., all models finetuned on ScholaWrite show an improved performance compared to their baselines. BERT and RoBERTa achieved the most improvement, while LLama-8B-Instruct showed a modest improvement after fine-tuning. As detailed in Appendix E, more training epochs and encoder-decoder architectures of BERT variants are assumed to be the reason for significant improvement compared to LLMs. This aligns with findings from Grasso and Locci (2024); Yu et al. (2023), which shows that RoBERTa and BERT can often match or even outperform LLMs for the text classification tasks. Those results demonstrate the effectiveness of our ScholaWrite dataset to align language models with writers’ intentions.
5.2 Iterative Self-Writing
Setups
During iterative self-writing (Figure 4), a fine-tuned model processes LaTeX-formatted seed document (as “before-text”) with a context prompt to predict the next intention, then revises the text (“after-text”) accordingly given prompt. The revised document then serves as the new seed for the next iteration. This process repeats until a set iteration limit (e.g., 100) is reached. All models use the same train (80%)-test (20%) split across experiments. See Appendix E and Figure 14 for training details.
We fine-tune Llama3.1-8B-Instruct (Llama-8B-SW) and compare it to vanilla Llama-8B-Instruct (Llama-8B-Zero) and GPT-4o. Also, seed documents were derived from LaTeX-formatted abstracts of four award-winning NLP papers on diverse topics Zeng et al. (2024); Lu et al. (2024b); Du et al. (2022a); Etxaniz et al. (2024) (Appendix LABEL:table:seed-entry-johnny-LABEL:table:seed-entry-latxa).
Metrics
We evaluated lexical diversity (unique-to-total token ratio), topic consistency (cosine similarity between seed and final output), and intention coverage (unique writing intentions used over 100 iterations). For human evaluation, three native English speakers with LaTeX expertise assessed outputs from Llama-8B-Zero and Llama-8B-SW on accuracy (alignment with predicted intention), alignment (similarity to human writing), fluency(grammatical correctness), coherence(logical structure), and relevance (connection to the seed document) - Refer to Appendix LABEL:sec:appendix:human-eval. Accuracy was judged per iteration, while other metrics used pairwise comparisons. Inter-annotator agreement (IAA)777The IAA scores are 0.84 (SW) and 0.76 (Zero) for accuracy, all 100% for alignment, fluency, coherence, and 49.8% (SW) and 100% (Zero) for relevance. was measured using Krippendorff’s alpha for accuracy and percentage agreement for others.
Metrics | Model | Seed 1 | Seed 2 | Seed 3 | Seed 4 |
Accuracy | SW | 21.0 | 10.3 | 18.3 | 15.3 |
Zero | 35.7 | 29.7 | 45.3 | 43.3 | |
Alignment | SW | 0 | 0 | 0 | 0 |
Zero | 3 | 3 | 3 | 3 | |
Fluency | SW | 0 | 0 | 0 | 0 |
Zero | 3 | 3 | 3 | 3 | |
Coherence | SW | 0 | 0 | 0 | 0 |
Zero | 3 | 3 | 3 | 3 | |
Relevance | SW | 1 | 3 | 2 | 1 |
Zero | 3 | 3 | 3 | 3 |
Results
Figure 5 shows that Llama-8B-SW consistently produced the most lexically diverse words, generated the most semantically aligned topics (Seeds 1 & 2), and covered the most writing intentions (except Seed 3). These results underscore the value of ScholaWrite in improving scholarly writing quality generated by language models.
However, our human evaluation (Table 5) revealed that Llama-8B-SW generated less human-like writing, in terms of fluency and logical claims. It also struggled with generating texts aligned with the predicted intentions. See Appendix Tables LABEL:table:human-eval-seed1 to LABEL:table:human-eval-seed4 for more details. Despite the weaknesses, Llama-8B-SW still produced more relevant content (Seed 2), which aligns with topic consistency trends in Figure 5, highlighting the usefulness of ScholaWrite dataset in certain contexts.
Moreover, Llama-8B-SW exhibited the most human-like writing activity patterns over time (Figure LABEL:fig:writing-step-intention-all-model), which frequently switches between implementation and revision and covers all three high-level processes. Llama-8B-Zero and GPT-4o tend to remain in a single high-level stage throughout all 100 iterations of self-writing. Compared to Appendix Figure 12, which depicts frequent transitions across all three stages in an early draft (e.g., the first 100 steps), Llama-8B-SW most closely replicates human writing behaviors in iterative writing tasks. These findings reinforce the potential of ScholaWrite in helping LLMs emulate human scholarly writing processes.
6 Conclusion and Future Work
We present ScholaWrite, a first-of-its-kind dataset capturing the end-to-end cognitive process of scholarly writing, comprising nearly 62K LaTeX keystrokes collected via our custom-built Chrome extension. This dataset, sourced from ten graduate students with varying levels of scientific writing expertise, is further enriched with expert annotations based on a novel taxonomy of cognitive writing intentions inspired by Flower and Hayes (1981); Koo et al. (2023). Through several experiments, ScholaWrite shows its value for advancing the cognitive capabilities of LLMs and developing cognitively-aligned writing assistants, enabling them to mimic the complex, non-linear, and intention-driven nature of human writing.
Limitations and Ethical Considerations
Appendix A More About Data Collection Process
A.1 Participant Recruitment & Demographics
A.2 Technical Details of System Implementations
When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel888To prevent privacy concerns, the extension filters out keystroke data from any unauthorized Overleaf projects. Please see Appendix A.2 for more details.. When each of these actions999Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page. occurs, the extension uses ‘diff_match_patch’ package101010https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 6). Then, the extension will send the array along with metadata (e.g., time stamp, author ID, etc.) to the backend server.
For any Overleaf project that consists of multiple LaTeX files, we also collected all keystrokes from subfiles associated with the main LaTeX file. Our comprehensive data collection process captures the end-to-end writing processes of the participating authors across all parts of Overleaf projects. This approach ensures that our dataset reflects the full scope of scholarly writing including edits made in auxiliary files such as files of each section, appendix, bibliography, etc.
Privacy Concerns
To prevent any issue of private data collection, we designed the backend of our Chrome extension to fetch only the IDs of the Overleaf projects that participants consented to share during the recruitment process and filter out participants’ keystroke data from any unauthorized projects. We used Google Sheet API to retrieve ID information that we collected during the recruitment process.
A.3 Data Post-Processing
[…]
Appendix B More About Writing Intention Annotations & Taxonomy
B.1 Annotator Recruitment
B.2 Detailed Annotation Process
Appendix C More About the Taxonomy of Scholarly Writing Process
During the planning stage, the writers engage in a process of generating and organizing raw ideas, arguments, or content structures that were not introduced in the previous trajectory. Based on the plan, the writers implement their plan by drafting full sentences and paragraphs and structuring the contents tangibly. At the same time, the writers enter the revision stage by improving the quality of their implemented sentences and LaTeX objects in terms of linguistic styles, format, or information accuracy. Particularly, spans of keystrokes whose intentions involved any changes but did not change the meaning of original texts are classified as Revision. For those edits that show changes in the meaning, we considered them as Implementation. Furthermore, if an author repeatedly adds, removes, and revises text back and forth until a sentence is completed, we consider this process as part of text production. Any subsequent changes made to the sentence after it is finished are considered revision. Table 1 presents the comprehensive, complete definitions of each intention of end-to-end scholarly writing process, identified from ScholaWrite dataset.
Appendix D ScholaWrite dataset statistics
Table 6 presents the overall statistics of our ScholaWrite dataset. Table 8 shows the distribution of intention labels per Overleaf project from ScholaWrite.